Isabelle has 3 nickels, 2 quarters, 2 dimes, and 7 pennies. What's an expression that can be used to find how much money Isabelle has altogether? How much money does Isabelle have?
Answer by ikleyn(52781)


Money = 2q + 2d + 3n + 7p = 2*25 + 2*10 + 3*5 + 7 = 50 + 20 + 15 + 7 = 92 cents.